OpenStudy (anonymous):

If you rationalize the denominator of 1/(7\sqrt(5)-3\sqrt(3)) then you will get (r)sqrt(5)+(s)sqrt(3)/n where r, s, and n are all positive integers (with no common factors).

OpenStudy (anonymous):

Find \(r,s,n\)?

OpenStudy (anonymous):

\[\frac{1}{7\sqrt5-3\sqrt3}\times\frac{7\sqrt5+3\sqrt3}{7\sqrt5+3\sqrt3}=\frac{7\sqrt5+3\sqrt3}{(7\sqrt5)^2-(3\sqrt3)^2}=\cdots\]